OpenStudy (anonymous):

need help finding derivative of f(t)= t^2/3 log subscript 9 (square root of t + 5 )

OpenStudy (anonymous):

\[f(t)=t^{\frac{2}{3}}\log_9(\sqrt{t+5})\]

OpenStudy (anonymous):

yes perfect !

OpenStudy (anonymous):

chain rule and product rule as well as of course the power rule

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

ok lets go slow \[(fg)'=f'g+g'f\] right?

OpenStudy (anonymous):

where to begin

OpenStudy (anonymous):

put \[f(t)=t^{\frac{2}{3}},f'(t)=\frac{2}{3}t^{-\frac{1}{3}}\] that part is easy yes?

OpenStudy (anonymous):

separate them and take derivative of first part. which is fairly eadsy

OpenStudy (anonymous):

but the log is scary

OpenStudy (anonymous):

the hard part is \[g(t)=\log_9(\sqrt{t+5})\]

OpenStudy (anonymous):

your in calculus...learn to love logs now...they're your best friend

OpenStudy (anonymous):

so first off \[\frac{d}{dt}\log_b(t)=\frac{1}{\ln(b)x}\]

OpenStudy (anonymous):

but you don't have \[\log_9(t)\] you have \[\log_9(\sqrt{t+5})\] so for this you need the chain rule

OpenStudy (anonymous):

therefore \[\frac{d}{dx}\log_9(\sqrt{x+5})=\frac{1}{\ln(9)\sqrt{x+5}})\times \frac{d}{dx}\sqrt{x+5}\]

OpenStudy (anonymous):

so is it \[\log _{9}(5+t)^{1/2} = (1)/(\ln(5+t)^{1/2}\]

OpenStudy (anonymous):

oops i did it wrong then. sorry. i tried to follow ur formula

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

hold on lets make it easy ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\log_9(\sqrt{x+5})=\frac{1}{2}\log_9(x+5)\] by that simple property of the log that exponents come out front as coefficients

OpenStudy (anonymous):

OHHHHH

OpenStudy (anonymous):

i hope this step is clear, i am using \[\log(x^n)=n\log(x)\]

OpenStudy (anonymous):

so i was on the right track. so is this the final derivative ?

OpenStudy (anonymous):

yes that makes more sense :)

OpenStudy (anonymous):

sorry but my time is running out... :-S

OpenStudy (anonymous):

this should make life much simpler. and in fact we should have started right away with \[f(t)=t^{\frac{2}{3}}\log_9(\sqrt{t+5})=\frac{1}{2}t^{\frac{3}{2}}\log_9(x+5)\]\]

OpenStudy (anonymous):

so is this the final asnwer ?

OpenStudy (anonymous):

or u just rewrote the original problem

OpenStudy (anonymous):

no i wrote the original problem. we go right to the answer now

OpenStudy (anonymous):

oh okay whew.

OpenStudy (anonymous):

\[\frac{1}{2}[\frac{2}{3}t^{-\frac{1}{3}}\log_9(x+5)+t^{\frac{3}{2}}\times \frac{1}{\ln(9)(x+5)}]\]

OpenStudy (anonymous):

pulled out the one half right at the beginning, then used the product rule

OpenStudy (anonymous):

so iam going to submit it. omg i hope its correct. thanks soooooooo much for putting up with me. i really appreciate it. you are a life saver

OpenStudy (anonymous):

its a really long asnwer.

OpenStudy (anonymous):

\[=[\frac{1}{3}t^{-\frac{1}{3}}\log_9(x+5)+t^{\frac{3}{2}}\times \frac{1}{2\ln(9)(x+5)}]\]

OpenStudy (anonymous):

either way

OpenStudy (anonymous):

oh distribute the 1/3 through? that would be more complicated. okay i will stick with wat u have

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

i just distributed the 1/2

OpenStudy (anonymous):

it is along answer because it is a product rule problem

OpenStudy (anonymous):

you gotta type this in good luck!

OpenStudy (anonymous):

ohhhhhhh okay. omg i mso nervous. i hope its right. thanks a lott through i really appreciate it .

OpenStudy (anonymous):

urgh got it wrong. :(